solve completely. explain your answer please

media%2Feec%2Feec64e86-a93d-45e8-9620-e8

solve completely. explain your answer please
Answer #1

a)

Let +x be towards leftward and +y be towards upward

X and Y components of Final Horizontal velocity

Vfx=VCos60

Vfy=-VSin60

Here final horizontal velocity is equal to initial horizontal velocity i.e

Vfx=Vox

Horizontal distance traveled

d=Vfxt

25 =(VCos60)(1.5)

V=33.33 m/s

Therefore

Vfx=Vox=33.33Cos60 =16.67 m/s

Vfy=-33.33Sin60=-28.87 m/s

From

Vfy=Voy-gt

-28.87 =Voy -9.81*1.5

Voy=-14.1525 m/s

Since

V_{ox}=V_{o}Cos\phi =16.67----------------1

V_{oy}=V_{o}SIn\phi =-14.1525------------2

2/1

tan\phi =\frac{-14.1525}{16.67}=>\phi =tan^{-1}(0.849)=-40.34^{o}

Therefore

From 1

Vo=16.67/Cos(-40.34) =21.87 m/s

From

Y=Yo+Voyt-(1/2)gt2

when the ball hits ground

0=h+(-14.1525)(1.5)-(1/2)(9.81)(1.5)2

h=32.27 m

b)

Vo=21.87 m/s

c)

\phi =-40.34^{o}

d)

Since angle is negative ,ball is thrown below horizontal